Trace the curve y²=(x+1)(x-1)² by showing all the properties you use to trace it.
1
Expert's answer
2019-10-01T10:29:24-0400

"y_1=(x-1)\sqrt{(x+1)}\\
y_2=-(x-1)\sqrt{x+1}"

"y_1'=\sqrt{(x+1)}+\frac{(x-1)}{2\sqrt{(x+1)}}=\frac{3x+1}{2\sqrt{(x+1)}}\\
y_1''=\frac{3}{2\sqrt{(x+1)}}-\frac{3x+1}{4(x+1)\sqrt{(x+1)}}=\\
\frac{3x+5}{4(x+1)\sqrt{(x+1)}}\\
y_2'=-\frac{3x+1}{2\sqrt{(x+1)}}\\
y_2''=-\frac{3x+5}{4(x+1)\sqrt{(x+1)}}"

zero of "y_1'=0\ and \ y_2'=0\,is \, x=-1/3"

Since "y_1''(-1/3)>0,\ y_1" has minimum at this point.

Since "y_2''(-1/3)<0,\ y_2\" has maximum at this point.

Since "0\le y^2 \ and \ 0\le(x-1)^2" ,it follows that "0\le(x+1)\ and\ -1\le x"

Intersections with axis:

"y=0, \ (x+1)(x-1)²=0,\ x=\pm1"

"x=0,\ y=\pm 1"